Assume that the 60-dB bandwidth of a receiver is 12 kHz and the 6-dB bandwidth is 7 kHz. What is the shape factor of the circuit?

1
Expert's answer
2021-11-21T17:26:27-0500

Shape factor, SF, is the ratio of two bandwidth:


"SF=\frac{BW_{60-dB}}{BW_{6-dB}}=\frac{12}{7}=1.714"